2/5 = 5/10n + 2????????????????


The second one is: 7/10 = 3/8k + 1????????????????

Answer by Alan3354(21583)

2/5 = 5/10n + 2
Do you want to solve for n? You didn't say.
Multiply thru (all terms) by 10 n
4n = 4 + 20n
0 = 4 + 16n
16n = -4
n = -1/4
---------
The second one is: 7/10 = 3/8k + 1
Multiply by 40k
28 = 15 + 40k
40k = 13
k = 13/40